OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sqrt{x ^{2}-9} }{ x }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{9\tan ^{2}\theta } }{ 3\sec \theta }*3\sec \theta \tan \theta \]

OpenStudy (anonymous):

then \[\int\limits_{}^{}3\tan ^{2}\theta \]

OpenStudy (raden):

then use the identity : tan^2 (θ) = sec^2 (θ) - 1 see the integral becomes: